3. Find the maximum volume of a right - circular cylinder that can be inscribed in a right-circular cone of
radius R = 3 and height H = 9.
1
Expert's answer
2021-09-15T02:58:11-0400


Let radius of cylinder be r and height be h as in the diagram above.

So in ∆ABC and ∆ADE;

"\frac{AB}{AD}=\frac{BC}{DE}"

="\frac{(H-h)}{H}=\frac{r}{R}"

"\implies r=\frac{(H-h)}{H}R"

Volume of cylinder=πr2h

="π*(\frac{(H-h)^{2}}{H^{2}})R^{2}*h"

V(h)="\frac{πR^{2}}{H^{2}}(H^{2}+h^{2}-2Hh)h"

="\frac{πR^{2}}{H^{2}}*(H^{2}h+h^{3}-2h^{2}H)"

"\frac{dV(h)}{dh}=0" for maximum volume

So we differentiate (H2h+h3-2h2H) with respect to h

=H2+3h2-4hH=0

3h2-3hH-hH+H2=0

3h(h-H)-H(h-H)=0

(h-H)(3h-H)=0

"\therefore" h="\frac{H}{3}"

V="\frac{πR^{2}}{H^{2}}*[(H^{2}*\frac{H}{3})+(\frac{H}{3})^{3}-2(\frac{H}{3})^{2}*H]"

="\frac{πR^{2}}{H^{2}}*[\frac{H^{3}}{3}+\frac{H^{3}}{27}-\frac{2H^{3}}{9}]"

="\frac{πR^{2}}{H^{2}}*\frac{4H^{3}}{27}"

=4πH"\frac{R^{2}}{27}"

Vmax="4πH\frac{R^{2}}{27}"

r=3, H=9

Vmax="4*π*9*\frac{3^{2}}{27}"

=37.699 cubic units